Reports 1-1 of 1 Clear search Modify search
que - 12:36 Wednesday 18 March 2020 (13663) Print this report
Drift of feedback signal in IP and F0 with temperature.

As shown in the first figure the temperature has drifted from 23.7 °C to 22.9 °C in 60 days. Such a large change has produced an increase in the feedback signals of the IP coils. Particularly uncomfortable is H2 actuation which reaches -18,500 counts. Currently the limit in the medm screen is set to ±20,000 counts.

I changed the limit of H2 actuation to 23,000 counts and left the other two coils in 20,000 counts.

The second figure shows the drift of the F0 actuation signal.  It is currently -18,800 counts. It will not sature but it's already high also.

Below there is information relevant to BS suspension related to temperature:

  • Requested temperature for Type-B suspensions: 7420.
  • BS was assembled, in average, at 26.7 °C.
  • Requested temperature of booth: 25 °C.
  • Current temperature of booth: 22.9 °C.
  • Difference with requested temperature: 2.1 °C.
Images attached to this report
Comments to this report:
tatsuki.washimi - 13:30 Wednesday 18 March 2020 (13665) Print this report

All GAS filters and temperature are always monitored in this page (need KAGRA's common pass)

By the way, it is quite difficult to satisfy the request from BS and SR2 simultaneously with current system, because they are in the same clean booth with small distance.
You should introduce something to control each chamber temperature directly. (e.g. this one)

Images attached to this comment
tatsuki.washimi - 13:45 Wednesday 18 March 2020 (13666) Print this report


fabian.arellano - 14:24 Wednesday 18 March 2020 (13667) Print this report

The requested temperature of the whole clean booth is 25 °C, which is a compromise between what it's thought to be suitable for BS (26.7 °C) and SR2 (24.6 °C).

The information about the heating jackets is highly appreciated and will be passed up in the chain of command.

Search Help